# Machine Unlearning for Uplink Interference Cancellation

### v1

@@ Abstract
Machine unlearning (MUL) is introduced as a means to achieve interference cancellation within artificial intelligence (AI)-enabled wireless systems. It is observed that interference cancellation with MUL demonstrates $30\%$ improvement in a classification task accuracy in the presence of a corrupted AI model. Accordingly, the necessity for instantaneous channel state information for existing interference source is eliminated and a corrupted latent space with interference noise is cleansed with MUL algorithm, achieving this without the necessity for either retraining or dataset cleansing. A Membership Interference Attack (MIA) served as a benchmark for assessing the efficacy of MUL in mitigating interference within a neural network model. The advantage of the MUL algorithm was determined by evaluating both the probability of interference and the quantity of samples requiring retraining. In a simple signal-to-noise ratio classification task, the comprehensive improvement across various test cases in terms of accuracy demonstrates that MUL exhibits extensive capabilities and limitations, particularly in native AI applications.
@@ II Preliminaries and Problem Construction
where the independent and uncorrelated frequency-selective fading channel is $\textbf{h}_{1}\in\mathbb{C}$ and $\textbf{n}\sim\mathbb{CN}(0,\sigma^{2})$ is i.i.d additive complex white Gaussian noise. In this case, SNR with unit power is following
@@
With the assumption of $|\textbf{h}_{1}|^{2}>|\textbf{h}_{k}|^{2}$, a successive interference cancellation (SIC) algorithm [12] can be applied to capture $U_{1}$ throughout $t_{0}$ to $t_{2}$. However, this requires the knowledge of amount of interference users and channel state information (CSI) on both $U_{1}$ and $U_{j\neq 1}$ and considering the extensive time interval, it is a costly application under fast fading channels. On the other hand, a MUL algorithm is capable to cancel the interference with only a non-interfered retain set.
@@ II-A Model and Dataset Construction
A supervised SNR classification application is under investigation as an example for MUL interference cancellation. The dataset $\mathcal{D}_{U}=\{\mathtt{x}^{(i)},\gamma^{(i)}|i\in\{1,2,\cdots,T\}\}$ where $\gamma^{(i)}\in\mathbb{R}^{1}$ is the SINR of the $i$th block and the feature map of the received modulated information is denoted as $\mathtt{x}\in\mathbb{R}^{2}$. 2-dimensional IQ feature maps captured with follows
@@
Each $A^{2}$ map batch is obtained with $\mathtt{x}_{s}^{2\times T}\rightarrow\mathtt{x}^{(2T/A^{2})\times A\times A}$ which corresponds to $i$-th block transmission. Structure of $\mathcal{D}=\{\mathcal{D}_{1}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \mathcal{D}_{2}\}$ is follows
$\mathcal{X}=\begin{cases}\mathtt{x}_{(1)}\subset\mathcal{D}_{1},&\text{Retain Set},\\ \mathtt{x}_{(N)}\subset\mathcal{D}_{2},&\text{Forget Set}\end{cases}$
@@ III-A Evaluation
A simple way to evaluate the performance of $\mathcal{M}$ is to utilize $\mathcal{D}\backslash d_{u}$ to regenerate $\mathcal{Q}^{\circ}$ such that $d_{U}$ has no impact in it. Accordingly, $\mathcal{Q}^{\circ}$ is denoted as re-learned retain (RR) model. In this case, $\mathcal{Q}^{\circ}$ can be interpreted as the ground truth model for $\mathcal{Q}^{\prime}$. Therefore, $\mathcal{M}$ results with $\mathcal{Q}^{\prime}=\mathcal{Q}^{\circ}$ in an ideal unlearning process and it is desired that $\mathcal{Q}^{\prime}$ perform as well as $\mathcal{Q}^{\circ}$ for the same sample space $\mathcal{X}$. Thence, $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ can be used in both $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ for a fair comparison.
For the interference cancellation, $d_{U}=\mathcal{D}_{2}$. Moreover, previously established $\mathcal{D}_{U}^{\prime}=U_{1}$ allows to conserve $\mathcal{D}_{U}^{\prime}=\mathcal{D}\backslash d_{U}$. In one point of view, the interference source of $U_{2}$ is considered as the undesired forget set.
@@ III-E Data Collection and Unlearning Cost
In this section, the interfered uplink signal collection and the unlearning cost for this dataset can be found in terms of the number of samples needed to be unlearned. To continue with the $N$ user case scenario, the amount of MUL request is $N-1$. It is assumed $\mathcal{D}_{2}<<\mathcal{D}_{1}$ so that the unlearning process does not change the size of $\mathcal{D}$. $U_{1}$ always exists on each time frame and the probability of the $i$-th user is active with ${\mathcal{P}_{i+1}}$. Due to the superimposed operation on interference, the number of users does not affect the total sample number $K$. Furthermore, the amount of sample needs to be training is $K-\sum_{i=1}^{M}\sum_{j=1}^{N}d_{U_{j}}^{(i)}$ for $M$ amount of model. Assuming the same amount of samples is collected within each uniform $t_{0}$ time frame, corresponding to $\frac{K}{t_{0}}$ samples for each time frame. Therefore, the probability of $\mathcal{D}$ construction with $N\geq 2$ users is $\prod_{i=1}^{N-1}{{\mathcal{P}_{i+1}}}$ in a time frame. The event of at least one interference exist within a model is denoted as $E_{I}$ and it’s probability is given as below:
@@
Assuming each time frame is independent from another, the probability of $m$ out of $M$ models having IUI by $N-1$ users can be shown with binomial distribution, where the probability mass function of $m=\bigcup^{N-1}_{i=1}U_{i+1}$ for $0\leq m\leq M$ is given below:
@@
Therefore, the total amount of unlearned samples in each submodel (first sum) of the each request (second sum) is the following
@@ IV Numerical Results
IV-A Learning and Unlearning
IQ feature map is taken as grey scale image with $1$ input channel in the first convolution layer, and it outputs $32$ channel by using $3\times 3$ kernel. With forward pass data flow, first fully connected layer produces $128$ unit with $26\times 26$ spatial size that following with $\mathrm{ReLU}$ activation function. Using the $128$ units from previous layer, the second fully connected layer obtains a $5$ dimensional logits corresponding to each class and a $\mathrm{Softmax}$ is applied the logits to obtain class probabilities. Learning is demonstrated with $0.001$ learning rate for $50$ iteration, optimizing with Adam optimizer. The classifier accuracy in each batch is the ratio of correct SNR estimations to the total batch number. The complete model can be found in 111 https://github.com/riguwen/MULforIC.git.

### v2

@@ Abstract
Machine unlearning (MUL) is introduced as a means to achieve interference cancellation within artificial intelligence (AI)-enabled wireless systems. It is observed that interference cancellation with MUL demonstrates $30\%$ improvement in a classification task accuracy in the presence of a corrupted AI model. Accordingly, the necessity for instantaneous channel state information for existing interference source is eliminated and a corrupted latent space with interference noise is cleansed with MUL algorithm, achieving this without the necessity for either retraining or dataset cleansing. A Membership Inference Attack (MIA) served as a benchmark for assessing the efficacy of MUL in mitigating interference within a neural network model. The advantage of the MUL algorithm was determined by evaluating both the probability of interference and the quantity of samples requiring retraining. In a simple signal-to-noise ratio classification task, the comprehensive improvement across various test cases in terms of accuracy demonstrates that MUL exhibits extensive capabilities and limitations, particularly in native AI applications.
@@ II Preliminaries and Problem Construction
where the independent and uncorrelated frequency-selective fading channel is $\textbf{h}_{1}\in\mathbb{C}$ and $\textbf{n}\sim\mathbb{CN}(0,\sigma^{2})$ is i.i.d additive complex white Gaussian noise. In this case, SNR of $U_{1}$ with unit power is following
@@
With the assumption of $|\textbf{h}_{1}|^{2}>|\textbf{h}_{k}|^{2}$, a successive interference cancellation (SIC) algorithm [12] can be applied to capture $U_{1}$ throughout $t_{0}$ to $t_{2}$. However, this requires the knowledge of amount of interference users and channel state information (CSI) on both $U_{1}$ and $U_{j\neq 1}$. Considering the extensive time interval, it is a costly application under fast fading channels. On the other hand, a MUL algorithm is capable to cancel the interference with only a non-interfered retain set.
@@ II-A Model and Dataset Construction
A supervised SNR classification application is under investigation as an example for MUL interference cancellation. The dataset $\mathcal{D}_{U}=\left\{\mathtt{x}^{(i)},\gamma^{(i)}\,\middle|\,i\in\left\{1,2,\ldots,T\right\}\right\}$ where $\gamma^{(i)}\in\mathbb{R}^{1}$ is the SINR of the $i$th block and the feature map of the received modulated information is denoted as $\mathtt{x}\in\mathbb{R}^{2}$. 2-dimensional IQ feature maps captured with follows
@@
[…]
$\mathcal{X}=\begin{cases}\mathtt{x}_{(1)}\subset\mathcal{D}_{1},&\text{Retain Set},\\ \mathtt{x}_{(N)}\subset\mathcal{D}_{2},&\text{Forget Set}\end{cases}$ (6)
@@ III-A Evaluation
A simple way to evaluate the performance of $\mathcal{M}$ is to utilize $\mathcal{D}\backslash d_{u}$ to regenerate $\mathcal{Q}^{\circ}$ such that $d_{U}$ has no impact in it. Accordingly, $\mathcal{Q}^{\circ}$ is denoted as re-learned retain (RR) model. In this case, $\mathcal{Q}^{\circ}$ can be interpreted as the ground truth model for $\mathcal{Q}^{\prime}$. Therefore, $\mathcal{M}$ results with $\mathcal{Q}^{\prime}=\mathcal{Q}^{\circ}$ in an ideal unlearning process and it is desired that $\mathcal{Q}^{\prime}$ perform as well as $\mathcal{Q}^{\circ}$ for the same sample space $\mathcal{X}$. Thence, $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ can be used in both $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ for a fair comparison. For the interference cancellation, $d_{U}=\mathcal{D}_{2}$. Moreover, previously established $\mathcal{D}_{U}^{\prime}=U_{1}$ allows to conserve $\mathcal{D}_{U}^{\prime}=\mathcal{D}\backslash d_{U}$. In one point of view, the interference source of $U_{2}$ is considered as the undesired forget set.
@@ III-E Data Collection and Unlearning Cost
In this section, the interfered uplink signal collection and the unlearning cost for this dataset can be found in terms of the number of samples needed to be unlearned. To continue with the $N$ user case scenario, the amount of MUL request is $N-1$. It is assumed $\mathcal{D}_{2}<<\mathcal{D}_{1}$ so that the unlearning process does not change the size of $\mathcal{D}$. $U_{1}$ always exists on each time frame and the probability of the $i$-th user is active with ${\mathcal{P}_{i+1}}$. Due to the superimposed operation on interference, the number of users does not affect the total sample number $K$. Furthermore, the amount of sample needs to be training is $K-\sum_{i=1}^{M}\sum_{j=1}^{N}d_{U_{j}}^{(i)}$ for $M$ amount of model. Assuming the same amount of samples is collected within each uniform $t_{0}$ time frame, corresponding to $\frac{K}{t_{0}}$ samples for each time frame. Therefore, the probability of $\mathcal{D}$ construction with $N\geq 2$ users is $\prod_{i=1}^{N-1}{{\mathcal{P}_{i+1}}}$ in a time frame. The event of at least one interference exist within a model is denoted as $E_{I}$ and it’s probability is given as below
@@
Assuming each time frame is independent from another, the probability of $m$ out of $M$ models having IUI by $N-1$ users can be shown with binomial distribution, where the probability mass function of $m=\bigcup^{N-1}_{i=1}U_{i+1}$ for $0\leq m\leq M$ is given below
@@
Therefore, the total amount of unlearned samples in each submodel (first sum) of the each request (second sum) is
@@ IV Numerical Results
IQ feature map is taken as grey scale image with $1$ input channel in the first convolution layer, and it outputs $32$ channel by using $3\times 3$ kernel. With forward pass data flow, first fully connected layer produces $128$ unit with $26\times 26$ spatial size that following with $\mathrm{ReLU}$ activation function. Using the $128$ units from previous layer, the second fully connected layer obtains a $5$ dimensional logits corresponding to each class and a $\mathrm{Softmax}$ is applied the logits to obtain class probabilities. Learning is demonstrated with $0.001$ learning rate for $50$ iteration, optimizing with Adam optimizer. The classifier accuracy in each batch is the ratio of correct SNR estimations to the total batch number. The complete model can be found in the soruce code 111 https://github.com/riguwen/MULforIC.git.